OpenStudy (anonymous):

Determine an equation of the line that is tangent to the graph of f(x) = sqrt (x+1) and parallel to x-6y + 4= 0

OpenStudy (anonymous):

I found the derivative , which is the slope is 1/ 2 sqrt (x+1) what do i do next?

OpenStudy (turingtest):

you need to find when that the derivative is the same as the slope of the given line what is the slope of the given line?

OpenStudy (lgbasallote):

get the derivative of sqrt (x-1)...that's your equation of the tangent

OpenStudy (turingtest):

what is the slope of x-6y + 4= 0 ?

OpenStudy (turingtest):

your derivative is wrong, by the way...

OpenStudy (anonymous):

is it.... cuz if it is parallel, i can just plug in that slope to the equation..

OpenStudy (anonymous):

how is it wrong...

OpenStudy (turingtest):

maybe it's just your notation\[\frac d{dx}\sqrt{x+1}=\frac d{dx}(x+1)^{1/2}=\frac12(x+1)^{-1/2}\]

OpenStudy (anonymous):

yea..it looks the same i have to solve it by limits

OpenStudy (turingtest):

we need when that is the same as the slope of the given line. m:\[\frac12(x+1)^{-1/2}=m\]oh you need the definition, eh? ok...

OpenStudy (turingtest):

\[\frac{\sqrt{x+1+h}-\sqrt{x+1}}h\]\[=\frac{\sqrt{x+1+h}-\sqrt{x+1}}h\cdot\frac{\sqrt{x+1+h}+\sqrt{x+1}}{\sqrt{x+1+h}+\sqrt{x+1}}\]\[=\frac{x+1+h-x-1}{h\sqrt{x+1+h}+\sqrt{x+1}}=\frac{h}{h\sqrt{x+1+h}+\sqrt{x+1}}\]\[=\frac{1}{\sqrt{x+1+h}+\sqrt{x+1}}\]which in the limit is\[\frac1{2\sqrt{1+x}}\]so what have you got for m ?

OpenStudy (anonymous):

umm..isnt the answer of the derivative = the slope?

OpenStudy (anonymous):

and for the line, i got 1/6 as the slope..it looks wrong...i dont know

OpenStudy (turingtest):

yes, and we want to know when that slope is the same as the slope of the given line\[\frac1{\sqrt{x+1}}=\frac16\]yes, 1/6 is what I got

OpenStudy (anonymous):

so do i find x using 1/2 sqrt (1+x) = 1/6 ?

OpenStudy (turingtest):

yes, you need to know the point x for the point-slope form when we make our line

OpenStudy (anonymous):

oh ok thanks!! i got x= 8 :)

OpenStudy (turingtest):

so now you use the calculus version of point-slope form:\[y-y_1=f'(x_1)(x-x_1)\]

OpenStudy (anonymous):

don't really get it ..

OpenStudy (anonymous):

do u find a point and then plug in to y2- y1 /x 2- x1 = m?? i got 6y - x + 3 = 0 which is different from the textbook ans